OpenStudy (anonymous):

given u = [1,2,3] and v=[2,-2,-6] use dot product to find a unit vector perpendicular to both

OpenStudy (anonymous):

\[ \mathbf{w}\cdot \mathbf{u} = 0 \\ \mathbf{w}\cdot \mathbf{v} = 0 \]First find \(\mathbf{w}\). Then make it a unit vector.

OpenStudy (anonymous):

Consider letting \(\mathbf{w} = [w_1,w_2,w_3]\). You'll get a system of equations.

OpenStudy (anonymous):

cos u could easily find it by cross product (: but then my teacher asks us to do by dot

OpenStudy (anonymous):

Well get it.

OpenStudy (anonymous):

and you could use any point for w1, w2, w3? how could you know which points to match with both vectors

OpenStudy (anonymous):

Do you know how to solve a system of equations?

OpenStudy (anonymous):

You have two equations and three variables.... that means you have one free variable. Just set the free variable to 0 to make your life easy.

OpenStudy (anonymous):

or set it to 1

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

Do you have to write out the equations for you?

OpenStudy (anonymous):

I never solve this by dot, only with cross (:

OpenStudy (anonymous):

\[\begin{array}{rrl} \mathbf{w}\cdot \mathbf{u} = 0 & \implies \begin{bmatrix} w_1 \\ w_2 \\ w_3 \end{bmatrix} \cdot \begin{bmatrix} 1 \\ 2 \\ 3 \end{bmatrix}=0 &\implies w_1+2w_2+3w_3=0 \\ \mathbf{w}\cdot \mathbf{v} = 0 &\implies \begin{bmatrix} w_1 \\ w_2 \\ w_3 \end{bmatrix} \cdot \begin{bmatrix} 2 \\ -2 \\ -6 \end{bmatrix}=0 &\implies 2w_1-2w_2-6w_3=0 \end{array} \]

OpenStudy (anonymous):

Now, we let \(w_3=1\) since it is a free variable. That gives us: \[ \begin{array}{rcl} w_1+2w_2+3w_3&=&0\\ 2w_1-2w_2-6w_3&=&0 \\ w_3 &=& 1 \end{array} \]

OpenStudy (anonymous):

@pinky_cute1995 Do you remember how to solve systems of equations?

OpenStudy (anonymous):

so w1 = -1 ?

OpenStudy (anonymous):

You should solve for \(w_1\) and \(w_2\) Do you know how to do substitution or elimination?

OpenStudy (anonymous):

okayyy

OpenStudy (anonymous):

I got it

OpenStudy (anonymous):

the vector is [1,-2,1]

OpenStudy (anonymous):

An axle has two wheels of radii 0.95 m and 0.15 m attached to it. A 30 N force is applied horizontally to the edge of the larger wheel and a similar weight hangs from the edge of the smaller wheel. a. the net torque acting on the axle b. the maximum weight that the 30N force can pull up

OpenStudy (anonymous):

You have to now make the vector into a unit vector.

OpenStudy (anonymous):

I know how to do the unit vector hehe

OpenStudy (anonymous):

@pinky_cute1995 mention me in the other question if you want me to respond to it.

OpenStudy (anonymous):

okayyyy

OpenStudy (anonymous):

Divide it by the magnitude...

OpenStudy (anonymous):

\[ ||\mathbf{w}|| = \sqrt{\mathbf{w}\cdot \mathbf{w}} = \sqrt{1^2+(-2)^2+1^2} = \sqrt{6} \]

OpenStudy (anonymous):

it's okay

OpenStudy (anonymous):

So the unit vector is \[ \left[\frac{1}{\sqrt{6}},\frac{-2}{\sqrt{6}},\frac{1}{\sqrt{6}}\right] \]

OpenStudy (anonymous):

I know how to do that (:

OpenStudy (anonymous):

omggg

OpenStudy (anonymous):

You said you didn't...

OpenStudy (anonymous):

read it again ^^

OpenStudy (anonymous):

Okay

OpenStudy (anonymous):

flutter

OpenStudy (anonymous):

FUUCK

OpenStudy (anonymous):

anyway thank you again

OpenStudy (anonymous):

you are my angel hahaha

OpenStudy (anonymous):

For the physics one

OpenStudy (anonymous):

did you draw the picture?

OpenStudy (anonymous):

which one ?

OpenStudy (anonymous):

the plane one or the axle

OpenStudy (anonymous):

the two wheels thing

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

I try

OpenStudy (anonymous):

but then it kinda weird :[

OpenStudy (anonymous):

Hmmm, well to get the net torque we need the value of the weight

OpenStudy (anonymous):

Ok we could answer in terms of the weight.

OpenStudy (anonymous):

If the mass is \(m\) then the net torque is going to be: \[ 30 \times 0.95 - m\times9.8\times 0.15 \]

OpenStudy (anonymous):

so what is toque ?

OpenStudy (anonymous):

Torque is the force vector cross product with the radius vector.

OpenStudy (anonymous):

The radius is vector is just the vector from the axis of symmetry to the point where the force is applied.

OpenStudy (anonymous):

In this case since the radius and vector are perpendicular, the magnitude of the cross product is the the product of the magnitudes...

OpenStudy (anonymous):

But you should read up on it.

OpenStudy (anonymous):

okay